Suppose you have a solution that contains 0.0470 M Ca2 and 0.0910 M Ag . If solid Na3PO4 is added to this mixture, which of the following phosphate species would precipitate out of solution first?

Ca3(PO4)2 is the answer to part 1.

When the second cation just starts to precipitate, what percentage of the first cation remains in solution?
Answer #1

Ca3(PO4)2 <-------->3Ca2+ + 2(PO4)3-
=>Ksp =[Ca2+]^3 x[PO43-]^2

Ksp of Ca3(PO4)2 is1.3 x 10-32

Ag3PO4 <----> 3Ag+ + PO43-
Ag3PO4 Ksp = 1.80 x 10-18

when does the second cation begins to precipitate

1.80x10-18 = [Ag+]3 [PO43-]
1.80x10-18 = [0.0910]3 [PO43-]
[PO43-] = 2.388x10-15
2.388x10-15 M the second begins to ppt,

but you asked: What is the concentration of the "first cation to precipitate" at that point:

1.3 x 10-32 = [Ca2+]3 x[PO43-]2

1.3 x 10-32 = [Ca2+]3 x[2.388x10-15]2


1.3 x 10-32 = [Ca2+]3 [5.70x10-30]

[Ca2+]3 = 0.00227 Molar

[Ca2+] = 0.13158 molar
======================================...

What percentage of the initial Ca2+ concentration remains in solution when the Ag+ begins to precipitate?

0.0470molar Ca2+ / 0.13158 = 0.357 =35.7%
Ca2+ remain in solution